Zjs log

From Yavuz Group
Jump to navigation Jump to search

--Zjsimmons (talk) 19:52, 19 March 2015 (CDT)

  • Reading a lot: papers, theses, etc. trying to understand the line broadening for our system. I really want to know the decay rate to be able to plug it into our simulations. More on that subject later.
  • Got another isolator so i'm able to start working with the wide-tune laser. So far i see a mode-hop free tuning range of about 5GHz, that should be sufficient for our purposes, but i really want to get it working better. Initially didn't seem to work very well, but i'm hopeful that increasing the load is helping give more repeatable/reliable piezo scan performance.


--Zjsimmons (talk) 11:50, 11 March 2015 (CDT) Quick Update:

  • Wide-tune laser is designed, drawn and built. Seeing initial laser feedback. Now to see if I can get it performing well and to see if i can optimize the tuning range.
  • Re: Laser linewidths: The interferometer has been set up for measuring linewidths at 1064, i think other Zach aligned it for us. I looked at two lasers we have on hand: the spare laser set up in our default way and the laser set up for rotating the grating about its face. Both had linewidths less than a MHz. Also looked at the experiment laser, initially the linewidth was quite large (few MHz), which may explain some difficulties with the crystal and why we haven't seen any narrow burned features. Attenuating the fast feedback gain to the diode and changing the laser driver seemed to bring the linewidth down to less than a MHz. When locked there was some evidence of further narrowing but the shape appears suspect and difficult to fit. We should probably look at this further.
  • Re: EIT and experimental simulations: we looked at the delay between green and fluorescence but it was difficult to conclude when the fluorescence 'turned on'. Figuring out the effective gamma to use in simulations is still an unresolved question. We were able to measure AOM switching and observed times of 100ns or so, as expected. Another question is what effective density to use, have an estimate but this needs to be looked at more carefully as it contains a lot of assumptions. I don't think we're going to have much hope of seeing things until we have parameters really nailed down and a decent handle on a simulation.
  • Re: In the crystal: trouble seeing anything when focused, but why?

--Zjsimmons (talk) 15:44, 2 March 2015 (CST) Recent work, three main irons in the fire:

  • wide-tune laser design: finally have Zemax working, need to figure out more exact cavity optical path length measurement to inform CAD laser design.
  • green beam on-off: installed switches, have the NI-6002 putting out TTL pulses, doesn't seem very stable so far, perhaps pursue a clock-linked design, maybe we could get better stability. Built a TTL or-gate circuit to allow switches to be controlled by delay box or 6002, so far it doesn't work properly.
  • EIT parameters: investigating with my simulation to try and decide on good parameters for seeing EIT in the experiment. So far I see evidence of EIT, but there are still a lot of parameters to play with.

--Zjsimmons (talk) 14:22, 16 February 2015 (CST) Last few days: Little progress, having trouble digesting the new information. We have been making progress but there is still so much we want to do. Did order some parts to hopefully build a laser with a larger mode-hop-free freq sweep and as a very simple NI DAQ to use to turn RF switches on and off.

Ideally we will have evidence of EIT before the end of the month. Also would love to have a wider absorption scan of the overall (inhomogeneously broadened) absorption feature showing things like a spectral hole and side/anti- holes.

Things to better understand:

  • How much power should we use? I think we need to know this if we want to know how long of EIT pulses to use as well.
  • Spectral hole width: this is determined by hole-burning power, i.e. more power=larger width, but what exactly is the relationship? Does this even matter that much when talking about pumping?
  • Fraction pumped back: When looking at transmission, it appears that a lot of population can be pumped back into the hole. However, this is somewhat misleading because we may have enough optical depth at that laser power to absorb our laser many times over in the crystal adjacent to that spectral hole in freq. Presumably, if we were to turn up the laser power of the probe scan to the point that it would be not totally absorbed, we would see that absolute contribution of the population pumped back from the other hyperfine ground states. Thus far, when turning up the probe power somewhat (perhaps a x10 factor) we don't see much of an intelligible change. I would expect the overall amount pumped back to appear smaller. Perhaps we just don't have enough probe sweep power, to see the difference.

Alternatively, we should look at the fluorescence. This could give us an observation that doesn't have the relative background. Fluorescence should relate directly to how much is there. We will i think want to understand this better for collecting data for our upcoming EIT experiments anyway.

Things to build:

  • ULE cavity.. this could really help with our laser drift..
  • Wide sweep laser, this would be great but will be some work.
  • Implement switches on RF sources so we can turn them on and off through a combination of the delay box and computer.



--Zjsimmons (talk) 17:47, 9 February 2015 (CST) OK, spent a little time today with the He. Some observations:

  • spectral hole width depends on burning power. Green power of a few uW is sufficient to create noticeable features in the absorption. But what about burning all the way through, etc. hard to see as you don't really get much fluorescence when you try and scan with low power..
  • We don't have our pump AOM in the right spot, we'll have to put it in after the control AOM in the control beam path if we want to get the correct spacing.
  • Overall population MAY be decreasing, this would make sense if population was accumulating in the other hyperfine state, we should get the other AOM set up to investigate this.

Lots of things to get sorted out. Also, thinking about building a laser better suited to wide scanning...


--Zjsimmons (talk) 18:46, 6 February 2015 (CST) Week in review. It's been an exciting week, we're finally doing some science! after spending a lot of time building things.

  • Mon-put the ref cavity in a can; i should probably get some windows for that to further isolate it from the environment. We also did a little work with the crystal.
  • Tue-more work with the crystal, hole-burning etc, temp in the lab spiked, destabilizing things.
  • Wed-temp was wacked up, did some reading on spectral hole burning.
  • Thur-Got back to the crystal. Spent a ton of time trying to get the laser int he exact right spot. The mode-hop-free tuning range is about the same size as the feature we want to see so it doesn't work very well for scanning, as you have to get it int he exact right spot and what you're looking for also has to be in the middle of the scan range. Remarkably, it cooperated for a little while. We were able to slowly ramp the laser piezo and lock the SHG cavity to the moving laser. This lets us scan the whole inhomogeneous linewidth and observe a burned feature 40MHz wide in the overall inhomogeneous linewidth. Needs to be improved if we are going get decent data, but we were able to see something!
  • Fri- step back to document, try and digest what we've been looking at, some notes with scope shots:

File:2 2 data summary.pdf

Note: the laser mode-hopping doesn't appear to cause a problem for the fiber amp, this is good news!

A few questions/issues:

  • Need a laser with a larger mode-hop-free tuning range if we are going to be able to scan over the whole thing.
  • Need to resolve why it appears we can pump so much back into our hole; need to reconcile abs and fluorescence observations and straighten that out.

--Zjsimmons (talk) 17:33, 30 January 2015 (CST) Experimenting with spectral hole burning. Burning with a 10mW laser, looking with a sweep that is ~1-2uW. Observations:

  • Spectral hole is perhaps wider than we expect: We need to look at the laser linewidth, but i was thinking we would we see a hole ~1MHz wide and instead its more like ~10-15MHz wide.
  • Temperature change-> ref cavity length-> laser wavelength->hole position: We heated the cavity with the heat gun for a fraction of a minute, see the transmission peak (spectral hole) fall out of the scan range. Also as the cavity length was changing and laser changing color, when we burn a hole the hole is large and shallow, this makes sens as the laser was effectively scanning.
  • Hole persists for a long time: We see the hole minutes after its creation, as it drifts out of and into view.
  • Big question: does burning a hole wipe out an adjacent hole?, no we can burn holes next to each other: two spectral holes

File:Two spectral holes.pdf

Also, i changed out the mirror for the invar ref cavity and it worked basically immediately, probably dumb of me to order that other cavity mirror. I wonder if that would have better stability performance? We should probably put it in a can and see.


--Zjsimmons (talk) 18:23, 29 January 2015 (CST) We are finally back to doing science again! In the last week or so we have the laser locking well to a cavity and seeding the fiber amplifier. Fiber amplifier provides way more power to the SHG cavity. There is still ringing on the SHG output, but it seems to work OK; we could potentially mess with the lockbox parameters to get a cleaner lock. Nick has the AOM beams set up.

Today we finally went looking for hole burning again. We see evidence of it, but it doesn't look as narrow as we would hope, appears to be more like ~5-10MHz wide, when it should be more like the laser ~1MHz (locked actually more like 50kHz?- need to measure this). Perhaps this cavity is fluctuating significantly, (few MHz) or there is some other issue. Presumably, we could up the burning power but burn for less time, giving us a narrower feature? So far we have not observed such behavior yet. So far we have looked with about 1uW of power and burned with much more, up to ~1mW or so for a second.

We're doing science again! Next:

  • look at inhomogeneous broadening again, might as well with the more stable laser.
  • mess with lock parameters to get more stable/narrower laser?
  • ref cavity change?



--Zjsimmons (talk) 17:30, 21 January 2015 (CST) Added a pick-off to monitor laser on spectrum analyzer/wavemeter. Seems well behaved compared to the previous laser. This is quite encouraging. Have the fiber amp seed launch moved toward to the inside of the table, should protect it a little/keep it out of the way. Next is to mode-match the fiber laser output to the SHG cavity.

Spent some time dialing in the seed, current parameters:

1055.073nm Laser Details
Settings: 64.79mA 22deg C 44.8V DC on laser piezo
Power Budget: 2.05mW seed .95mW to lock ~125uW to monitor launch

--Zjsimmons (talk) 13:20, 20 January 2015 (CST) Swapped out the diode laser for the new one, it seems to be working well. Is much more efficient, 50mA of current provides more output power than 180mA to the previous laser. Lock seems to be working well as well, but it's hard to tell if that's the laser or just that there's a little more power going to the cavity.

Hooked up and seeded the fiber amplifier, with it quite in the lab, the laser and amp stayed locked for an hour no problem. Output power seemed basically constant, to within a percent or or ability to measure.

Question: so if the laser drops lock will that mess up the amplifier? -checked this, if the lock drops, the seed is still there, just not locked so we're ok.

Do need to add a pick-off in the system for the wavemeter/spectrum analyzer.


--Zjsimmons (talk) 13:33, 19 January 2015 (CST) Came in on Saturday and built a new lock-box circuit with the new board. Just re-hooked it up and it seemed to work! First time i've seen locking via PDH.

Couple of circuit issues to fix:

  • need to move the switch that grounds the ramp BEFORE the input resistor to that op-amp
  • jumpers that i though would be clever are really unnecessary/wrong, a bnc short at the input does the job.
  • put piezo fb switch after gain section so we can observe the gain change the signal magnitude
  • note: error signal DC offset pot direction changes with invert switch: this is not ideal..

--Zjsimmons (talk) 11:59, 15 January 2015 (CST) So it appears that the mirror reflectivity for the invar cavities is so high it is making them hard to work with. Ordered a broadband (lower reflectivity) mirror from thorlabs, we'll see if that helps.

In the meantime, i've been working with the other "new" laser i have been working on, i would like to get it in the bullpen. There was really no reason it wasn't working, i just needed to keep playing with it until i found some feedback. Interestingly, i think the tilt of the grating had insufficient travel to get feedback at the correct wavelength, apparently this was why it wasn't working/showing any feedback.

Initially i noticed feedback with the polarization of the laser diode oriented 90 deg to what I expected, this lead me to compare to the other diodes but they are in fact the same. I then changed the grating angle to get the correct output wavelength. Maybe i was just unable to get the grating angle close enough before to observe any feedback.

Initially the feedback was not very good, it was not a single peak (in output power vs tilt space), but a hill with many sharp peaks that have different output wavelengths, each local peak about a 20%increase in output power. Eventually i was able to get better performance by messing with tilt and collimation as well as grating angle. I adjusted the asphere position to try and get better feedback but that suffered from the fact that just pushing on the lens and changing its tilt in the mount changed how well the feedback worked; this is a pretty significant problem with our lasers, relying on adjusting collimation by turning the lens in its mount. I followed up by carefully measured the grating angle and tried to get it closer to what it's supposed to be. At some point, while moving the grating to a closer angle to what i expect, the feedback performance greatly improved. Now see about a factor of 6 due to optical feedback, .25mW to >1.5mW output power at about 22mA input current.

For a 1200 rulings/mm grating and 1055nm output, i'd expect about 39.3deg off the normal, but ended up at about 37.7. This is not unreasonable as i'm not able to measure the angle super accurately. Looking at the other laser design, that turns about the grating face, that one appears to have an angle of about 41 deg. Interestingly, these seem a lot more angled than our other 1055nm laser.

Lesson learned re: tuning up lasers, if the grating angle isn't right to get feedback, you're not going to see anything. Useful to make sure you're at the right angle (or close) if you want to have hope of finding feedback.


--Zjsimmons (talk) 19:53, 12 January 2015 (CST) Crunch time. So i machined 2 invar reference cavities, completed before the break. David is using one and i'm working with the other. So Far I have not been able to get good mode-matching, not good enough coupling to lock to it. Potential issues: -higher finesse than the other cavity so the ramp has to be better? -perhaps we need input wedge to avoid disturbing the seed laser?

Also, ordered new circuit boards for the lock box, tried to incorporate most everything i could think of. Omitted the op-amp based ffb and generally streamlined it. Hopefully it works better.

Getting locked to a ref cavity is our number one priority.


--Zjsimmons (talk) 15:59, 17 December 2014 (CST) So feedback electronics issues are driving me crazy and i'm giving that a rest. Serendipitously, the material for the invar fabry-perot cavities just arrived and i've started working on that. Hopefully after doing this, looking at the circuit with fresh eyes will yield some ideas/progress.


--Zjsimmons (talk) 18:56, 12 December 2014 (CST) Network analyzer is working better. So far haven't been able to get everything working, but at least we have an additional tool at our disposal. Lots of things to try.


--Zjsimmons (talk) 20:46, 11 December 2014 (CST) Re: Network Analyzer, ok, this thing is driving me crazy. If you go to too low of a stop freq, the structure disappears, although it doesn't give you an error or anything. This appears to be because if you scrutinize low frequency you need to decrease the resolution bandwidth to compensate. i.e. i think you need to look more accurately, hold the frequency more steady, longer. This means that the scan will take longer; perhaps my impatience got in the way. In any case, if you set it to auto, it should take care of getting an adequately small bandwidth given your other scan parameters. Note: you can also reduce the number of points to make the scan take less time.


--Zjsimmons (talk) 01:26, 11 December 2014 (CST) for lack of other ideas i've been putting the lockbox circuit in multisim to verify that the gain and phase-margin are sufficient for our application. No startling revelations as I looked at it when building the first box. A few ideas: -50ohm coupling resistor? -insufficient gain? Check error signal magnitude on the SHG lock and see if they're vastly different.. -something just not connected? since the ramp and the FFB work, there's a pretty narrow amount of stuff that may be messed up, look in the circuit? -evaluate the circuit with a network analyzer?


--Zjsimmons (talk) 12:46, 9 December 2014 (CST) Another day, another crack at circuit debugging.. -tried putting inverter before the lockbox, that doesn't seem to fix it. -invert fast feedback? putting inverter right before the fast feedback messes up the laser, linewidth looks like crap. -upped the gain in the slow feedback branch by x100, this makes it the same as the hexagon group lockbox schematic, still not seeing any slow locking at higher gain. -boo, perhaps i should simulate the circuit to look at bode plots/examine the circuit with a network analyzer?


--Zjsimmons (talk) 16:01, 8 December 2014 (CST) Re: circuits One issue is coming into focus: the source of the 10s of kHz noise on the circuit rails, that seems to bleed through into the error signal. It appears that its origin is the heater in the lm399s. I'm going to try running them without the heaters and see what happens. Running without the heaters squelches the kHz level noise and doesn't seem to prevent any ffb locking. Maybe this is the way we should run them.

I tried an inline inverter. Tough to use as the input to the driver cannot go negative and the lock box doesn't really put out a negative DC. As a result, when you flip (the positive) signal you get something negative or close, not much room to operate. Also, one of the inverters seems to rail which is problematic.

Also, interestingly, looking at the previous lockbox, there doesn't seem to be any discernible square-wave bleed-through to the error signal monitor. Seems to be i had the input set up wrong when i converted the differential amp to one input. Instead of grounding the other opamp input, i was actually looking at the differential of the power supply rails, that where the square-wave noise bleed through was coming from. That's taken care of, but the slow feedback still doesn't work. hmm. gains? try inversion again?

--Zjsimmons (talk) 19:07, 4 December 2014 (CST) Have not been able to get the slow feedback working. Hmm.. Tried a variety of things: -Scrutinized the switching -Tried locking with slow feedback by itself inverted, didn't see any evidence of locking. -I know the LM399s seem to be mysterious noise generators. Tried putting some caps across their heaters, as i've seen this in some circuit drawings, didn't seem to help with error signal issue, not sure if it helped with rail noise.

Perhpas: -circuit power supply could be better? -try inverting slow branch -too much/little gain - change cap and resistor values in fb branch -bad component? but the ramp works fine, most all of that is working -some strange coupling? it is weird we can see discontinuities on the error signal that correspond to the square pulse, do we see those in the other lock box?

--Zjsimmons (talk) 19:36, 3 December 2014 (CST) Made a little bit of headway today with the PDH.

  • Tuned up the FP cavity modes; ended up moving the cavity and mode-matching lens a fair amount, probably about an inch, putting the distance from the front mirror very close to where ABCD predicts it should be. Lesson learned: turns out it is much easier to move the lens than the cavity. Was able to get it pretty much single-mode at times.
  • Strung together the PDH electronics and it worked! well we could see an error signal anyway. Error signal looks pretty much like the Black paper. Did not have to struggle with the phase-shift in order to get a proper-looking error signal.
  • Stuck the lock-box in the set-up. We can turn on the fast feedback and see the central resonance broaden, as we expect, as we zoom in on the peak (i.e. adjust the piezo DC and shrink the ramp), the fast feedback intermittently locks on its own. This is pretty encouraging, so at least the FFB seems to be working. However, when we turn from ramp to lock, it doesn't lock. There is a problem with the slow (piezo) feedback branch. Probably the first thing to look at is the switch.

--Zjsimmons (talk) 22:13, 2 December 2014 (CST) Spent much of the day messing with the Fabry-Perot cavity. I removed the cavity and checked to see what the beam looks like in the position of the cavity; it looks very similar to what my ABCD matrix script predicts if i remove the input coupler, so i am confident that my mode-matching is decent. I mentioned before that i did not see a Finesse above about 20, it turns out these were not the cavity resonances we were looking for. Nick suggested looking at more extreme orientations. I turned the cavity 90deg in its mount, i.e. rotated the whole can, and then searched for more peaks. Then i saw much sharper peaks, finesse>500. My leading theory is that when i glued the piezo back on, i must have glued it off-center. This led to the cavity resonance mode being at a weird angle and thus hard to find. Perhaps the crappy peaks we were seeing were etaloning between a cavity mirror and the windows? or some other cavity-like effect? Those peaks were very clean, we didn't see neighboring modes, they also benefited slightly from turning the polarization, two differences from what we see now.

Now the finesse looks like it's more like 500, 1.5GHz FSR cavity so about ~few MHz wide resonances, comparable linewidth to the laser. Also the cavity features are narrow enough that we can resolve the EOM sidebands. I stuck a cube in the beam path and now i can resolve corresponding dips in the reflection off the front of the cavity. Spent time tuning up the cavity mode-matching but was unable to kill a feature sitting on the side of the main peak.

Hopefully we can get the electronics all set up now that this cavity is working better.

--Zjsimmons (talk) 17:46, 1 December 2014 (CST) Yikes, it's Dec, where has the semester gone. I have the "1030" Fabry-Perot that we have from QC set up. It doesn't look that great, but hopefully works well enough for us to set up the PDH electronics. Planning to work on that tomorrow.

In the process of ordering mirrors and Invar to build 2 cavities that would be same dimensions, hopefully very similar to our ULE reference cavity. A spare may be of use to David on his experiment.

Re: 1055 laser, boo i can't seem to get any feedback from the set-up that i would have thought would have worked the best. I'm using the laser diodes from the order F325732 that Jared acquired. Number 2 diode, i was only able to get perhaps 20% feedback and there were definitely 2 lobes in the output mode for some collimations. As a result, i switched to the other diode, number 3 and it also appears to have structure on the output mode, less symmetrical lobes, but there is definitely a second spot you can see for some collimations. Now i can't seem to get any feedback. I've played with the collimation, grating angle, grating tilt(of course) cavity length, cleaned the diode and lens(the lens is kind of crappy), tried another lens, and still haven't seen any feedback, very strange. Perhaps the diodes are crappy, but i would sure like to get it working and see something working well before using our new 1055 wavelength-selected diode. We'll see, i could break down and try one of our new diodes.

--Zjsimmons (talk) 13:18, 28 November 2014 (CST) A few updates. I spent much of last week modifying/constructing a lockbox for the PDH set-up utilizing the ULE cavity. I feel like i was able to apply some lessons learned from the last time around to aid assembly this time, principally putting it in a larger box made it easier to put together. Unfortunately, the cavity has a fatal problem and will be returned. They mirror glass is oriented the wrong direction giving structure on transmission and acting like a grating. Hopefully in the meantime we can get something else working to at least let us set up some of the electronics. Toward that end, we're working on two fronts:

a) we're trying out a "1030nm" cavity borrowed from quantum computing. The cavity is about 101mm in length and contains an internal label: HHR-R1030 R=100mm 1030nm 99.9805%. This gives it a FSR of about 1.5GHz.

I could not see very good peaks with it earlier this week but it appears that the laser was at fault. I am seeing decent peaks now, although they are not great. Based on mirror reflectivity we would expect very high Finesse, >1000, more like 15000, using labled reflectivity:.999805. However this is for 1030nm, at 1055 it could be significantly lower. However even if it was only 99% (couldn't find a datasheet/know exactly what the mirrors are) at 1055, we'd still expect a finesse of ~300. I observe a finesse of more like 20. Could there be something wrong with the laser? 80MHz seems crazy wide for our laser linewidth. Or is the cavity reflectivity really only <90% at 1055nm? Or is there something i'm missing. Need to spend some more time messing with this cavity.

b) I'm planning to build a temporary stop-gap cavity out of invar or super-invar with the same dimensions as our ULE cavity. I figure this would hopefully allow us to move forward while minimizing time lost as we don't really know how long it will take to get the ULE cavity replaced.

--Zjsimmons (talk) 17:03, 13 November 2014 (CST) This week has not been so productive in the lab.

  • updated the poster for MCAW, ended up taking way longer than i would have liked but perhaps that reflects that we're making progress and needed to thoroughly update!
  • ordering a bunch of misc opto-mechanics as well as RF gear for the PDH set-up. I'm really liking this surplus store BMI surplus; they have lots of stuff.
  • still haven't been able to get good feedback out of my new laser with the old diode, perhaps it's the diode, the output mode does look a little strange..

--Zjsimmons (talk) 18:11, 5 November 2014 (CST) I moved the SHG cavity, and launched into it with the revised laser and TA. Picked Nick's brain about mode-matching and had to refresh my memory on that. Wrote a quick script to get my bearings. Nick suggested having telescope lens on translation stage: that way that lens would help with beam size and the coupling lens close to the cavity would help with waist position. After messing with the lens and coupling tune-up, saw nice locking. I'm not sure why it worked well, but tuning up the cavity seemed to make it better. Perhaps the laser is better behaved and that is making a difference. Note: interestingly you can see noise (e.g. your voice) on the lock. Perhaps we should invest in sound insulation. Note2: unfortunately, for whatever reason, it seems like when the lock drops sometimes it rails the piezo; not ideal behavior.

--Zjsimmons (talk) 12:55, 3 November 2014 (CST) OK, so last week i spent a lot of time building another laser. This one is basically the same as our go-to design, except the back of the mirror mount, adapter ring, and base-plate are all replaced by a single piece of aluminum. The collimation tube remains, but otherwise everything is one piece. The rational is that poor thermal conduction between the pieces in the former design may have been leading (in part) to poorer feedback and performance. Hopefully we will be able to evaluate the temperature performance somewhat with Eli's temp monitor project. Anyway, i just got the adjustment screws for it today so am getting ready to finish assembly and put it into service.

Otherwise, somewhat encouraging news with the previous revised laser design. I see good feedback. I fiddled with the collimation tube (not the collimation) but the position of the tube, and started to see good feedback. Interestingly, even thought the laser is centered at 1064 or so, i was able to easily get it to laser at 1055 just by adjusting the grating angle. This is encouraging. It also appears that the monolithic design of that laser has better temp control properties. I was able to dial in the wavelength by changing the temp; i haven't had much luck with this with our default laser design.

Note: I'm planning to write up all the laser findings as a lab white-paper. This should also provide some good spring-boards to student projects.

--Zjsimmons (talk) 18:00, 23 October 2014 (CDT) The next order of business is getting the laser to behave better. This should allow us to better see spectral hole burning and just make our lives easier as well as be confident we won't drop the seed to the fiber amplifier.

To that end, i spent some time in the last week putting together a new laser prototype. Perhaps this is an unworthy side project, but i am rationalizing it by a) i want to understand these issues in depth myself and b) it should provide a test-bed for our future students to try some investigations: including feedback as fn of collimation, decoupling grating rotation from translation, etc. I intend to document my findings as well as i think that our progress/documentation on this issue as a group is pretty thin and wold like to see it expanded upon/brought to a more satisfying (not conclusion necessarily) but point anyway.

So far it seems to control diode temp quite well, i.e. the bare diode emission moves about 1nm/3deg C, which is apparently reasonable. This diode is centered at about 1066, so i don't think we'll be able to pull it to get it where we want it, but i'll see what i can learn from it. Feedback doesn't seem to be working very well yet, i.e. I cannot see a sharp feedback feature, although i see some feedback.

In other news, we're baking out the can for the reference cavity and it seems to be working, the pressure has come down to about 5E-6 over the last couple days at about 100degF.

--Zjsimmons (talk) 18:33, 15 October 2014 (CDT) Recap of last 2 weeks: Spent a lot of time last week building parts for our reference cavity. Learned about Autocad as well as how to hold things from the center on the lathe and how to use the boring tool on the milling machine for making large circles/portions of circles. Interesting, but took longer than i would have hoped.

We're finally starting to see some effects in the crystal. Observed absorption. Just now starting to hopefully see evidence of spectral hole burning. See notes:

https://wiki.physics.wisc.edu/yavuz/images/9/9f/Potential_spectral_hole_burning.pdf

--Zjsimmons (talk) 16:56, 2 October 2014 (CDT) Recap of the last few days. I ordered a bunch of stuff: some 4W RF amplifiers for higher power SiO2 based AOMs, an 80MHz AOM, an 85MHz high-power surplus AOM similar to what we already have picked up, a 210MHz AOM for our pump beam, some 1055nm specific laser diodes, lots of gear. Have built some misc stuff in the last few days: a box for storing the PCFs, am helping get Eric started with that stuff. A mount to adapt an AOM to a mirror mount. A plexi box to protect the output fiber of the fiber amplifier.

As far as actual science, i set up a double-passed 85MHz AOM on the green laser. This is what we intend to scan to try and observe spectral hole-burning. Nick is working on the RF source control via GPIB. Hopefully we can observe a nice narrow feature.

9-25-2014, Nick and I took some absorption data; i have it in my notebook. Removing the lens definitely helped.

--Zjsimmons (talk) 17:49, 24 September 2014 (CDT) ok, just got back from a visit/vacation in SD. Getting back to business. Nick took the focusing lens out of the system, we can definitely see absorption qualitatively now, very cool, the lower intensity/less susceptibility to motion did the trick. However, the laser is not cooperating. It will not stay stable enough to let us scan through the resonance with the piezo. We have tried tuning up the feedback, collimation, etc, but so far it's not cooperating. very frustrating.


--Zjsimmons (talk) 21:40, 14 September 2014 (CDT) Yikes; it's been a month since i've added anything to this document. Very disappointing to see what i am currently struggling with is what i was struggling with the last time i made an entry. I am trying to take a crystal absorption scan but cannot get the 1055 seed laser to cooperate. I have tried messing with the collimation as i did last time but so far have been unable to get it to cooperate. Perhaps i just need to fiddle with it more, or perhaps the laser needs to be changed. Arg.

Work has been progressing however. I've been working on a 6 level density matrix simulation, want to simulate what happens in the crystal. Have it up and running and will probably generate and post a report about that pretty soon.

Also, have been helping get Chandler started building a laser and Eric started with the PCF stuff. Things are getting done, but the laser is really being a pain in the butt.


--Zjsimmons (talk) 00:32, 26 July 2014 (CDT) What did we learn today? Well today was a pretty prime example of why i was a bit anxious about building up stuff this late in my grad school run. I spent the day trying to get the 1055 seed laser diode to cooperate. It's capable of doing what we want but has been difficult, i did learn/re-learn a few things:

  • Tried different polarization orientations. We want the spatial elliptical spot to be horizontal, but changing the angle changes how well the light is reflected from the grating and and as a result, how well the optical feedback works. I could observe different orientations having different amounts of feedback (or absence of feedback) but it was not an obvious relation- probably in part because the method to turn the tube is kind of crude.
  • Tried changing the cavity length. This didn't seem to do too much.
  • Tried adjusting the collimation. This ended up being the most important adjustment. Initially i adjusted the collimation just down the table, but this was too short a distance to look. Then Jared helped me collimate it on the wall and the feedback performance upon tune-up was much better. Saw threshold test go from 200uW to 1.6mW, pretty dramatic. My suspicion is that if the output is better collimated, the feedback light reflected back through the lens focuses much better. If it focuses better, you get better feedback and more single-mode behavior. When it was tuned up, we were able to scan over several GHz just by adjusting the seed diode laser piezo. The SHG cavity even stayed locked while changing the wavelength.
  • Immediately after tuning everything up, the laser was much better behaved, we were able to scan over the abs feature without the laser messing up during hte scam.


--Zjsimmons (talk) 10:52, 24 July 2014 (CDT) Re: 1055nm Laser: So at group meeting Deniz suggested we look at the output from the diode without feedback and see where the emission was centered. It's actually around 1067nm and we are unable to get that output band down to 1055 by temp tuning. Running it as cold as we could we only see about 1062. This would explain in part why the laser was a bear to tune, we're operating it far from it's actual 'natural' wavelength. Looking at the datasheets for 1060 diodes we have, it appears the one Jared was using was the closest. We looked at that on the spectrum analyzer and it is actually pretty much right where we want it. 1055 falls between two of the features in it's free-running emission spectrum. Next i switched our two diodes. Note: apparently diode wiring colors conventions are not always followed. I had a scary moment mis-wiring the diodes because the chosen colors on one of the diodes were very strange. For future reference i'm going to check the connector whenever re-wiring. A couple observations:

  • Jared's 1060 diode operates at about 2x the current for some reason.
  • The diode seems less tunable via temperature. Perhaps this is related to operating at higher current.



--Zjsimmons (talk) 17:48, 22 July 2014 (CDT) Re: SHG Locking stability and laser tuning. It seems like we're making a little progress understanding what affects the shape of the error signal and positions of the cavity modes due to the different polarizations.

  • When the different polarization modes are piled on top of each other, we get a crappy error signal and our lock is not very stable. This can happen when we change the cavity length somehow or change the temperature which also changes the cavity lengths for the two polarizations.
  • The error signal does not have to look perfect/textbook, i.e. the modes don't have to be very well separated but they do have to be distinct. It appears it's not necessary to stress getting the modes well separated.
  • There are wiggles at 100Hz or so on top of the transmitted peak (it's really more of a band because of the fast feedback). These wiggles cause locking performance to suffer as well. This is difficult to untangle from the other issues. It appears laser feedback/current can be adjusted to suppress these wiggles. Then the transmitted power band looks smooth and the locked output is very stable.
  • As per Nick the temp bandwidth is a couple degrees so we can adjust it by a 1/2 a degree or so to modify the shape of the error signal and get better locking performance. Considering the tuning sensitivity to temp (16deg/nm), we would want to adjust the temp a bit on opposite ends of a 1nm scan for example.
  • It seems our number one priority now is squashing the wiggles on the transmitted band. Although difficult to achieve, when the wiggles are gone and the laser is locked, the output is very quiet. It's not clear what is causing those wiggles. I though it was the laser piezo drifting, but shorting it did not eliminate the problem.



--Zjsimmons (talk) 17:51, 19 July 2014 (CDT) Week in review: I've basically been working on 2 things, trying to get better lock performance and starting to write up the EIT eqns so we can look at that in the context of the green transition in Eu:YSO. Nick and I tried to look for absorption on thur but didn't see very good results; stability of the output power/lock issues are really fighting us. We did learn some things though:

  • tuning up alignment to get even photodiode response seemed to help the lock performance, perhaps it helps with noise to have the optical feedback balanced?
  • the quality of the laser makes a big difference to the lock performance. i.e. the laser may look single mode on the spectrum analyzer and the wavemeter reading may be very stable but touching the current or feedback makes the error signal look much better. When it's poor it may appear that a band of noise periodically passes through the peaks, disturbing everything. zooming in on the peak, it looks noisy.

Today i made some changes to the electronics.

  • Added 50 ohm impedance matching resistors to the photodiode inputs. I had been hesitant to do this on account of it shrinking our input signal but decided to try it on account of noise; perhaps the low impedance would be less susceptible from noise coming from the photodiodes or noise from some ground loop through the photodiodes. Indeed the error signal shrank somewhat, but was still plenty big enough to do feedback.
  • Added a new passive fast feedback branch that goes directly to the laser diode. I talked to Larry yesterday about our issues trying to feedback through the laser driver. He said that that's problematic for same two reasons we had been suspecting: ground/noise issues and the fact the laser driver mod input has very high gain, so you need very low noise. He suggested a fast feedback branch directly to the diode and sketched out a little circuit. Turns out we had a lot of little discarded fast feedback boxes lying around so i cannibalized one of those used the layout suggested by Larry. His circuit is nice and simple and passive, so we also don't have to worry about powering it. I stuck it inline immediately before the laser. As for an error signal pick-off, i added one inside the lockbox, picking off inside the fast feedback branch, immediately after the first op-amp buffer. This makes it so the phase is correct, and not inverted relative to the slow feedback branch. I didn't disconnect the existing fast feedback but left it in place. We could try sending that directly to he diode as well. Below is a PDF of the circuit.
File:Passive fast feedback.pdf
It consists of two sections. There is a filter on the main laser driver line: two 820uH inductors as well as a 10uF cap and 3ohm resistor to ground. This is to prevent fast feedback from going back into the laser driver. Also, the low pass may clean up the laser driver output. The second section is the fast feedback branch. There are low pass and high pass branches, so the gain is actually less in the middle of the frequency range. From Larry: apparently you want to suppress the gain a tad in the middle as that's where the laser will naturally respond. The high frequency gain is to fix phase response, so that there is not too much phase accumulation at high freq. Also, he recommended the whole circuit being preceded by 10dB attenuator. I ended up putting in 16dB as that seemed to work well. It's interesting how you really don't need any gain in the fast feedback branch.
  • The lock was much better when utilizing this fast feedback branch, the transmitted IR was very flat in output intensity, fluctuating less than 10%. Lock is a little more finicky, i.e. photodiodes seem to get unbalanced more frequently, this probably due in part to the 50inputs. Hopefully we're finally getting a handle on these issues. Also, this fast feedback branch should be applicable for locking the laser to the reference cavity.

--Zjsimmons (talk) 18:33 Made some changes today, see qualitative improvement in green lock. So we've been having issues with the locking performance, the transmitted IR and also generated green were intensity modulated at about 300Hz, a strange freq to see problems as it was not the cavity piezo resonance freq of about 420Hz. The SHG derived green looked at times almost like a pulse train as a result of the modulation. This was tracked to in part a suspected ground loop between driver and locking electronics. Unfortunately the fast feedback was needed to get good output power, but would also cause the modulation.

  • Swapped the cavity piezo for the new shorter travel (6um) one from Noliac. Glued everything together with epoxy so as to cut down on load weight. Somehow, realigning the cavity didn't turn into a huge project. Note: when you see some green, put the spots on top of each other and you're most of the way toward good alignment.
  • Performed the resonance test on the piezo/mirror system. We now see resonance at about 700Hz. This is higher but not dramatically so. Also, the resonance was qualitatively not as noticeable, so perhaps it is less strong as well which would be cool.
Upon locking with the new piezo, observed very robust stable lock, you can really rap on the table and it won't drop. Piezo/mirror system definitely less susceptible to noise, banging, disturbing the lock. However we still see the modulation present via the fast feedback. Turning the FFB on and tuning to the best gain the output looks pretty good. Would estimate fluctuation in about the top 25% of the transmitted IR or so. Turning the piezo gain all the way up saw no piezo resonance with FFB off. Green is relatively constant, i.e. not intermittent, although there is fluctuation in the top 1/2 or 1/3 so.
  • Swapped R30 in the circuit from 1k to 100k, this should boost the piezo feedback by a factor of 100. Since we saw no resonance at max piezo gain with the new piezo, i decided to change the resistor to give us more gain to play with. Now at higher piezo gain (adjusted via pot) we can observe piezo resonance. FFB squelches this but we still see some wiggling on the transmitted IR and Green. Now the wiggles are more like 100-150Hz. There also seems to be a tad more fluctuation in general at low piezo gain now, this looks like it could be at 60Hz, so maybe we are starting to see power supply AC noise and harmonics or something like that getting through. I dialed it down a bit and changed the 100k to 30k, you can still observe wiggles.
  • Next priority: find some way to squelch the FFB noise problem. Looking online, and from Deniz's comments, it seems that the laser driver modulation input is very susceptible to noise/ground loop issues. Hopefully we can do something. I put a 15k resistor in the BNC shield between the lockbox and laser driver; that didn't seem to help much if at all. Will do some brainstorming on it. Fortunately, i think the current performance is sufficient for doing a green absorption experiment.


--Zjsimmons (talk) 21:54, 11 July 2014 (CDT) Week in review: Frustratingly slow week; didn't seem to make much headway/i don't know anything. Re: Cowan Code: After trying about a bunch of different routes to get it working, i think i've more or less decided on the original prescription, the Alex Kramida pre-compiled version for PC. Despite the big incentive to pursue the un-compiled versions so as to be able to use the scripts, i haven't been able to get it to work easily. Here's what doesn't work, there's basically 2 routes that i worked on for a while(too long):

  • Compiling the code in Ubuntu didn't work as I couldn't get the right compiler. Both compilers mentioned in the documentation are a bit of an issue. Unfortunately the Intel Fortran compiler doesn't work with the new version of Ubuntu. It is possible that it would work with an older version of Ubuntu or a different Linux distribution. g77 is not readily available/very old so also a problem. The gfortran compiler also doesn't work with the code, at least according to the documentation.
  • Compiling the program in windows via cygwin also didn't seem to work either. The Fortran compiler (ifort) although available for free for windows is not available for windows. Besides, you can't run the make files in cygwin and have it utilize the windows ifort anyway. However, you can have it use the gfortran, the open source fortran compiler. Although it seemed to compile, the code didn't work. According to the documentation, apparently that compiler won't work.

Re: Laser linewidth issue: We finally made a little progress on this problem. So it appears there is a problem with the fast feedback, some sort of ground issue. Deniz showed pretty dramatically that just touching the BNC ground to the laser feedback terminal on the driver totally messed up the error signal. As much as i would like the electronics to not be responsible for problems, it appears that they are. I kind of figured.

  • If we tune up the gains, piezo gain basically low as workable, fast feedback not too low but not too high(where it looks best when ramping), we get a reasonably stable lock. If we turn the fast feedback gain too low we see wiggles at the piezo resonance freq in the transmission. If the fast feedback gain is too high we see slow oscillations at around 300Hz, evidently the noise on the laser that was responsible for the more or less pulsed behavior of our green light.
  • The Fabry-Perot apparently led us astray. Either we were looking at the peaks while the fast feedback was connected (i could swear it wasn't) or the cavity linewidth is not as good as we thought. I thought that the linewidth changed slightly when we tuned up the fast feedback by doing the threshold test, indicating not being limited by the cavity linewidth, but maybe that was wishful thinking and we were actually limited by the cavity linewidth.
  • Then again, that's why we wanted to look with the interferometer, which worked very well. With the fast feedback disconnected, we observed a lindewidth of about .5MHz, very reasonable. However, with the fast feedback connected, even if it's off, the linewidth is more like 5-10MHz. Even just touching the ground of the feedback connection BNC connector ground totally disturbs it.
  • One piece of information that doesn't jive is the frequency of the noise oscillations, ~300Hz? it's not a line freq or harmonic. One idea: perhaps there is some sort of feedback or cross-talk between the two feedback branches. What if the piezo feedback is modulating the cavity transmission a bit, i.e. gain at the SHG piezo resonance freq is not sufficiently low, there's some wiggle. If this modulation is being fed back some to the laser, maybe that's responsible to the increase in lindewidth. I don't know if something along those lines could be the mechanism, but if it were, swapping the piezo could really help. That's our next idea priority.

--Zjsimmons (talk) 16:45, 7 July 2014 (CDT)

Trying to nail down why the locked laser is misbehaving. The symptoms:

  • The lock is stable, but the transmitted IR laser power is not constant but modulated about about 200Hz with a modulation depth of about 50%. The The error signal has a saw-tooth characteristic. This causes significant problems to the output green power as the SHG basically exaggerates the problem and we see pulsed green output instead of continuous green. We have observed the problem go away but we don't know what caused it to disappear.
  • Additionally, when scanning the DC offset, the shape of the transmitted resonance changes it noticeably broadens depending on the DC offset. This may or may not be related to the modulation problem.
  • When zooming in on a peak with the DC offset and ramp size, as you get very close, the peak looks crappy, there is a lot of structure present which at least part of is at the piezo resonance frequency.
  • Tried replacing the lock-box ramp with an external ramp, the appearance of the peak when you zoom in is perhaps slightly cleaner, but about the same.
  • Tried turning the waveplate on the cube that separates the polarizations for the two photodiode channels and then adjusting the error signal with the lock-box. This did not seem to help or change anything.
  • Perhaps it's a cavity alignment issue? Could going to transmission mode help?

--Zjsimmons (talk) 20:01, 3 July 2014 (CDT) Touched some science this week!

  • Spent a lot of time messing with the Cowan code. A couple issues to work out but i think it is workable, i'll write it up eventually, once i get it figured out. Here's the short story: There is a version of the code by Alex Kramida that is compiled for windows, it includes the 4 core programs: RCN, RCN2, RCG, RCE. This runs from DOS but the output is not very easy to interpret. There are scripts that automate running the above programs and parse the output into something a little more useful. However these only work with Unix/Linux, although there is a workaround for in windows using a sort of Linux terminal emulator called Cygwin. This should work although i haven't been able to get it to work so far. Alternatively you can compile and install the programs in Linux, but i haven't been able to get that to work either. The third option is to rather than use the pre-compiled versions, to compile the programs via Cygwin in Windows. I haven't been able to get this to work yet, but i haven't spent much time trying either. In any case, one of these options should work and allow the use of the scripts.
  • The science: we put some green light through the Eu:YSO crystal at 5K. We're currently using the set-up with the tapered amplifier rather than the 1055 fiber laser as we're hunting around for the transitions and don't want to disrupt the fiber laser seed. We see some evidence of the transitions, not so much absorption but fluorescence; it looks pretty bright orange.
  • Interestingly the orange appears at only narrow green wavelength, but we get a lot more fluorescence when the crystal is warm
  • Transitions appear to be right where they are supposed to be from the literature. We expect it at 527.5381nm at see very bright fluorescence at 1055.0775/2=527.53875nm. Also see some orange in the vicinity of the adjacent site's peak at 527.6426nm.
  • Wavemeter and spectrum analyzer give different peaks. 1055.0818 on the wavemeter is 1055.004 on the optical spectrum analyzer. OSA seems to be off by about 0.08nm.
  • If you pull out the transfer line when the cryostat is cold, it warms up pretty slowly, might be able to use to take measurements as fn of temp
  • There is also evidence of absorption but it's not very good at this point. The green transmission is more like a a pulse train than a constant output, this needs to be fixed/addressed. SHG seems to make variations in cavity transmission much worse; i suppose it sort of squares them. Hopefully we can get the output more constant.


--Zjsimmons (talk) 13:14, 23 June 2014 (CDT) Holy cow can't believe it's been a couple weeks since i've updated my log. A few things have happened:

  • DAMOP. Great having it in Madison, got to see some old friends including Herman Batelaan and Jeff Sherman whom i hadn't seen since Seattle. Sort of surprising to me that i basically saw no other work using doped crystals, however made a good contact with Morgan Hedges. He worked under Matthew Sellers on some rare-earth doped crystal stuff. Been in touch with them and currently looking at some of their papers. I think this could be some very good folks to know int he field, folks who have some experience with these crystals that we're just starting to experiment with.
  • received the lockbox PCB from advanced circuits and wired it up, this took much of last week. seems to be working ok, although we're planning to change the piezo and box the SHG cavity to hopefully get more stable performance.
  • received the 1055 fiber amp from Nufern for the green laser, started setting it up.
  • just today i finished my documentation for the circuit and for building circuit boards, will post on the wiki. Building circuit boards is under Procedures. The lock-box documentation is under Equipment List, perhaps documentation should have it's own page, but this will work fine for now.


--Zjsimmons (talk) 21:33, 30 May 2014 (CDT) Week in review. Was out of town early part of the week for Noah and Teresa's wedding. Did get a couple things done this week:

  • ordered PCB for locking circuit, finally. Should be here in a week or so. The whole project was a good subject to learn about. I plan to document the method in a document and post to the wiki.
  • Nick and I updated our poster for DAMOP
  • Started work on the Eu:YSO crystal holder for the cryostat. Will finish machining next week.
  • Plugged in and turned on the Menlo Orange laser. Measured power of 110mW. According to email with Menlo, Chris McRaven, he measured 129mW. I don't know why there is a discrepancy. Could be they saw lower power upon adjustment of pulse length.
  • Note: last fri i finally submitted my paper review of the Optics Express paper on an Atomic Bragg Reflector. Really glad to have gotten that done. It really ate up more time than it should have but was a good experience to have done a paper review. Learned I would like to work on/understand the Stark effect better.


--Zjsimmons (talk) 16:11, 20 May 2014 (CDT)

Trying to get fast feedback to laser diode in place to see if it helps with cavity locking. So far it hasn't worked/i haven't been able to get good slow feedback locking. Difficult to untangle the different knobs we have control over. Issues/brainstorming to consider:

  1. Does polarization instability screw us?
  2. Can we trade a little bit of power (second waveplate, polarization incident on cavity) for better error signal performance?
  3. Tune up cavity peaks (IR) with crystal temp de-tuned and then zero in on correct temp? My suspicion is that adjacent peaks from different orders are messing up the error signal which is why it maybe isn't working well. If we tune up the IR, maybe this would be better to look at than trying to look at the green.



--Zjsimmons (talk) 18:00, 19 May 2014 (CDT) We observed some decent locking of the SHG cavity today. At highest, we saw intermittent power of about 9mW w/ about 65mW of 1055 incident, which is a pretty decent conversion ratio of ~14 or so. Observations/issues

  • Piezo is oscillating which is lowering output power. When the error signal gain is turned down, we see higher intermittent power but the lock is not very stable and drops. With higher gain, the oscillation is very pronounced.
  • Also put in place the fast feedback circuit to the laser driver, does not seem to help, in fact turning up the gain kills the lock. There could be many reasons for this. I am wondering if the phase is not correct. I would think it would be the same as the slow feedback branch and two inverting amplifiers shouldn't mess it up, but i would like to check this. Perhaps there is an issue with using long BNC and/or i would think that the bread-boarded circuit would be more susceptible to crappy performance at fast feedback frequencies. I do want to get this straightened out soon so i can order the board and get the permanent circuit built.
  • Also, it's becoming increasingly clear we should use a different piezo, something with higher resonant frequency would be very nice/easier to work with. I don't think we need the travel that the current piezo stack provides anyway.


--Zjsimmons (talk) 18:45, 16 May 2014 (CDT) Not much going on in the lab the last few days as i've been working on a report for my microscopy class. Tried looking at improving the resolution of HELM/SIM using a couple schemes. It appears they probably will not work, but it was interesting and i learned some stuff. Perhaps there's something there that could be developed further.

  • Locking circuit PCB layout is more or less ready to go. Want to test the circuit some more and make sure it works with the crystal in the cavity before getting it made. Also want to test the fast feedback. Nick has got the crystal in the cavity and we can see some single-pass green but it is definitely going to be more challenging to lock it. There's a lot more structure on the error signal and the cavity modes seem to jump around quite a bit.


--Zjsimmons (talk) 14:14, 13 May 2014 (CDT) Have seen a fragile lock:) Despite Multisim showing gain should be -20 dB at 400 Hz, it still seems prone to 420Hz (resonance) oscillation. Issues:

  • Switch to put bypass resistor to kill integrator when not locked seems important. Otherwise integrator will chug away and rail at some point. I'm also thinking having a low noise op amp for that stage should help prevent gain on op amp error for example. -added the switch.
  • Should we use a different piezo with a higher resonance frequency?
  • Piezo DC offset should only be positive- I changed this.
  • Lock seems ok. Also added a 100nF cap on the output/input to the sum with DC offset as a passive low pass filter. I'm sure this could be improved but i think we should try it for now and see what happens.


--Zjsimmons (talk) 22:51, 12 May 2014 (CDT) Spent today trying different locking gain stage configurations, tried adding an additional stage to increase gain. Discovered error signal was not being properly added to DC offset, this seems to be corrected but two problems have cropped up:

  1. Integral gain has a relatively large DC offset, i guess this is not to be unexpected, but sort of fights the goal of zeroing in on the resonance point by adjusting the DC offset of the ramp and then turning on the gain, i.e. killing the ramp and adding the gained error signal to ramp DC offset. This can be suppressed but comes at the expense of gain.
  2. Piezo is resonating. Despite lowering the gain and lowering the roll-off point to very low frequency, the piezo is now oscillating. Can this be eliminated? Perhaps we need a lighter mirror set-up to raise the mechanical resonance frequency? It may be a good idea to get more piezos anyway, in case of failure.


--Zjsimmons (talk) 18:44, 9 May 2014 (CDT)

  • Cannot seem to get the SHG cavity tuned up. Very frustrating, stuff in between my peaks is only getting worse, not smaller. Polarization from fiber is also drifting, annoying.
  • Trying to see locking. Not working so far.
  • Saw some strange behavior with TA, output power seems to have decreased somewhat since we first used it. With 5mW seed power:
Current (mA) ASE (mW) output (mW)
1000 16.6 120
1501 126 377
2001 274 632

Compared to before (4-14-2014), where:

Current (mA) ASE (mW) output (mW)
1005 18.9 165
1501 165 430
2002 - 838
2074 434 1W

This seems to be down a fair amount from initial output and spec, degradation? are we doing something wrong?


  • Shout-out to Jared, need to have power meter at 1064 not 780 when looking at 1064:P
  • Note: weird power meter behavior. Old power meter head seems to behave a little different than the newer one. The old one gives about a 20% higher reading between 1.06um and 1.05um. The newer one gives about the same reading for both wavelengths, something is fishy with that. Perhaps this explains some of the discrepancy between the TA performance now and before, if we measured before at 1.06 and now at 1.05, still a bit different. We should keep an eye on this.

--Zjsimmons (talk) 18:25, 8 May 2014 (CDT)

  • examined Josh's lockbox. It seems to short the positive rail momentarily when first connected to power and then recover. Very strange behavior. Occurs w/ ICs removed as well, so it doesn't seem to be the chips. Checked voltage ref pots, doesn't seem to be for example shorted diodes. curious.
  • Having trouble tuning up the SHG cavity, could not get very nice clean modes, always some garbage in between peaks. Subsequently spun the fiber launch and totally misaligned the whole shebang. I guess it will be good practice tuning it up. Still unable to get as clean of 'troughs' between peaks, may have to pick Nick's brain about it.

--Zjsimmons (talk) 17:30, 7 May 2014 (CDT)

  • Tried repairing the fault finder today. Turns out there are some optics inside, two elements in small brass rings, one attached to the laser. Looks like they were glued together at one time and had since broken loose. This was why we could not get light through fibers with the fault finder, even though it was on, it was not correctly launching into fibers. I aligned the two elements as best i could using the transmission through a multi-mode fiber and glued it back together. It sort of works, coupling to single mode fiber is not that great, but with a little fussing(rotating the fault finder and pulling the fiber out slightly) you can get a decent amount of light through a single mode fiber. We should probably get a new fault finder as I don't know how reliable this repair will be.
  • re: mysterious polarization noise, could it be we are blocking the beam with our fingers/fiber and that's why the power goes to zero?
  • Nick showed me how to tune up the SHG cavity today. It didn't seem too bad to mess with; it'll be good to not be afraid to adjust it.

--Zjsimmons (talk) 23:23, 6 May 2014 (CDT)

  • Figured out that the flat polished fibers were the cause of power fluctuations after the TA in 1055nm set-up. Using an angle polished fiber fixed the issue.
  • Power fluctuations are also not the reason the circuit is not locking the SHG cavity. Perhaps the gain roll-off needs to be at lower freq? as we are seeing some wiggles around 420Hz (piezo resonance freq). Maybe this is involved in failure to lock. We could also try adding fast feedback branch and see what happens.

4-6-2014 How about a log? Maybe this would work better that the paper lab notebook. Lab notebook seems ill-suited to many functions and i find myself not looking forward to using it, perhaps this will work better. Should be a more natural medium for computer-generated material. It's nice that it would be searchable as well. It could also host PDFs and reports and other documents. I kind of wish we would have started a wiki a while ago! -z